# Estimating LLM Consistency: A User Baseline vs Surrogate Metrics

### v2

@@ 4 Results
Here we first summarize the user study data collection (4.1), the results of which will be used to establish our baseline for consistency. To compare previous consistency methods to our user-rating based method, we follow a bottom up approach. First we compare how methods of estimating similarity between sentences (or LLM responses in our case) compare to user ratings (4.2). Next, we use these similarity scores (including user-based scores) to build consistency scores for individual responses (4.3). Here we also compare logit-based uncertainty estimation methods to our user baseline. Finally we aggregate consistency scores to a set of responses to obtain a prompt-level consistency score (4.4). For each step of the hierarchy, we compare automated methods of estimating consistency to our user-rating based ground truth. We find that while some methods are closer, none are able to match the user baseline. We further find that ensembling logit-based scores approximate the best-performing sampling method.
@@ 4.1 User Study Overview and Demographics
We recruited 2,976 participants from Prolific, asking each to rate five unique response pairs from different prompts. About 52% of participants were female and 47% male. Participants’ ages ranged from 18 to 88 with various education, ethnicity, and income backgrounds. More details are provided in table 5. We collected Prolific ID for compensating participants who completes our survey. We removed Prolific ID before doing any analysis.
@@
We used 100 prompts and generated 10 responses for each prompt. Each of the 10 responses per prompts then made up 45 response pairs. Removing duplicate response-pairs (including pairs with identical responses, see section A.2), we were left with 756 unique response-pairs, each of which were rated by five or more participants.
Overall, the prompts were on average 13.33 tokens in length and the responses 52.80 tokens long. We break down the lengths for prompts drawn from the two datasets and responses generated by the three models in table 2.
@@ 4.2 Response-Pair Similarity
Our investigation starts with comparing user ratings of the similarity between response-pairs to semantic similarity metrics used in prior work. Semantic similarity metrics is a core component of sampling-based consistency metrics. We call this comparison “response-pair similarities”.
@@
We evaluate the semantic similarities between each pair of responses using previously published methods and computed the correlation with human ratings. Comparing each of the four metrics evaluating the similarity of response-pairs—Bert, BLEU, Rouge, and USE Zhang et al. (2020); Papineni et al. (2001); Lin (2004); Cer et al. (2018)—to participants’ ratings, we found that no single score correlates best with how participants rate sentences’ similarities across all datasets and models. Notably, USE have the highest correlation coefficient with human ratings over all response-pairs ($\rho=0.75$) and for Llama while Bert performed the best for Mistral.
Interestingly, we found the correlation between participants’ ratings and five scores from previous work are higher for prompts from the CoQA dataset than the LMSYS dataset. Participants found response-pairs answering prompts from the CoQA dataset to be more similar to each other than response-pairs for LMSYS, by one level in the 6-point scale. To the best of our knowledge, Bert, BLEU, Rouge, and USE have not been used to evaluate LLM consistency with the more open-ended LMSYS dataset, which could have contributed to their lower correlation with human ratings.
Result 1: Among four existing metrics for evaluating the consistency of LLM responses based on similarity between pairs of sentences, USE performs the best in most cases. Additionally, we found that existing metrics better approximate human ratings on CoQA than on LMSYS.
@@ 4.3 Response to Response-Set Consistency
In section 4.2 we evaluted methods for measuring the similarity between pairs of responses. Consistency metrics, however, typically involve computing a score for each response that represents the extent to which that response is similar or different from other responses to the same prompt. Here we compare how a human-rating based consistency metrics performs in relation to those build with response-pair similarity metrics, as well as one additional metric that requires sampling multiple responses—SE Kuhn et al. (2023). We call this “response to response-set consistency”.
@@
Motivated by reducing the cost of generating multiple responses to a prompt for consistency estimation, we further explore if logit-based consistency estimations methods (which have negligible cost compared to response generation) can match human-based consistency scores. For this investigation, we used 16 different logits based scores—mean, minimum, maximum, and sum for each of DLR, Entropy, Probability, and LogProbability (explained in section 3.3)—and found none of them individually approximates the consistency of LLMs as well as the USE score (see table 4).
@@
Figure 2: At response to response-set level, our ensembling of 16 logit-based scores is as close of an approximation of human ratings as USE.
@@
As described in section 3.3, we used an ensemble of the 16 logit-based scores to attempt to approximate human ratings, with much better success. To determine which combination of the 16 logit-based scores can create the best ensemble, we used the Sequential Feature Selector (SFS) scikit-learn developers (2025). We ran SFS 1,600 times where each number of logit-based scores (between 1 and 16) is ran 100 times. Within each of the 1,600 runs, we used 10-fold cross validation and recorded the performance (i.e., Spearman $\rho$ and MSE) of our ensemble compared to human ratings. We found using all 16 logit-based scores resulted in the highest Spearman correlation coefficient and lowest mean squared error when compared to human ratings (see figure 5). We found our ensembling method with 16 logit-based scores had a higher correlation with human ratings than Bert, BLEU, Rouge, and SE. Our ensemble, when compared to the human ratings, performed as good as USE with a 0.002 better MSE (figure 2). Same as in section 4.2, we found at response to response-set level, existign metrics correlates with human ratings better on the CoQA dataset than LMSYS figure 6.
@@
Result 2: Using an ensemble of 16 logit-based scores can produce a consistency estimate as good as existing metrics at approximating human ratings. In addition to matching USE, our method benefits from not requiring the generation of multiple responses.
@@ 4.4 Prompt Level Consistency
As seen in the correlation difference between datasets table 3 and prior work Kuhn et al. (2023); Balabanov and Linander (2024) the prompt choice is a major contributor to consistency.
For example, responses to, “Where was the first modern Olympic Game?” is likely more consistent then responses to, “How can LLMs help users in their daily life?” As such, we next investigate consistency for a given prompt. Specifically, we aggregate response-level consistency from section 4.3 to obtain a single consistency score per prompt.
Figure 3: At per prompt level, our ensembling method and USE has the highest Spearman correlation coefficient with human evaluation of model-prompt consistency among existing metrics.
@@
Using prior defintion from section 3.3, we calculate the consistency score of each prompt using participants’ ratings and using metrics from previous work. We found that USE-based consistency scores best correlate with human-based consistency across all prompts (figure 3). We further note the consistency scores on the CoQA dataset correlates better to human ratings than those on LMSYS. This is similar to the bias we observed with pair-wise comparisons in section 4.2.
Ensembling of logit-based scores
Our ensemble of logit-based scores described in sections 3.3 and 4.3 performs as well as the USE when aggregated to the prompt level in approximating human ratings of LLM consistency. While it did not outperform USE, our method benefits from not needing to generate a set of responses and compare between responses to determine the consistency of an LLM at the prompt level.
We further investigate the number of responses needed to train an ensembled model with logit-based scores to predict the consistency of LLM. Due to the relatively small number of prompts and responses from our sample, we were not able to conclude how many responses are needed for such model to accurately predict LLMs’ consistency.
Result 3: Our ensembling method using logit-based scores performs as well as existing metrics in approximating human ratings of prompt level consistency. This approach enables estimation of LLM consistency without the need to generate multiple responses to a prompt.
@@ 5 Conclusion and Discussion
Using the human evaluation data obtained via our user study and subsequent experiments, we show that existing (automated) formulations of the consistency of LLMs are meaningfully different from our human judgment baseline. The best response to response-set consistency method achieves Spearman’s $\rho=0.80$, indicating correlation but not representation. Further, prior work has suggested that correlations considered “strong” in less precise contexts Schober et al. (2018) are not as meaningful for measuring success on NLP tasks Deutsch (2022); Bavaresco et al. (2024); Shen et al. (2023). This result holds regardless of whether we examine consistency at the per-response level (section 4.3) or at the prompt level (section 4.4).
Though sampling-based methods come closest to the human baseline, we show logit-based methods can be ensembled to approximate this performance ($\rho_{ensemble}=0.82$ vs $\rho_{\text{\emph{USE}}}=0.81$), creating an opportunity to avoid the sampling overhead.
We also find (section 4.2) that the difference in human judgement based uncertainty vs prior work is greater with real-world prompt datasets (LMSYS) compared to artificial ones (CoQA). Determining why this is the case is out of scope for our work, but we speculate that the research community’s focus thus far on artificial testing datasets might be a contributing factor.
@@
The discrepancy between the human baseline and existing methods raises important questions: How do imperfect consistency estimation methods affect downstream tasks? How would end-users be affected if shown consistency metrics Kapoor et al. (2024)? How are models affected when consistency measurements are part of the training cycle Liu et al. (2024)? The answers are unclear, but we believe they are worth investigating.
Finally, we urge researchers to consider the human evaluation baseline in future consistency methods revisions. We further recommend they (also) use real-world prompts for evaluation.
@@
Since our investigation requires logit-based scores which is often not accessible with black-box models, we used open-weights models. Such usage limits the generalizbility of our findings to black-box models. Additionally, the prompts and responses in our sample are relatively short in length (see table 2) despite us using realistic prompts from users Zheng et al. (2023), therefore, our findings may not generalize to prompts and responses of all length. Lastly, user studies are expensive to conduct. While we recuirted almost three thousand users, we were only able to evaluate 100 prompts with 10 repsonses per prompt. As we alluded to in section 5, more data from users are needed to establish a robust baseline for measuring LLM consistency.
@@
Our study is approved by our institution’s ethics review board. For the user study, we first provided an informed consent form to participants explaning the purpose of our survey, expected length, risks and benefits, as well as compensation. Participants who give consent to participate will proceed to the survey. As explained in section 3.1, responses shown to participants within the survey were filtered by the authors to remove any potentially harmful (e.g., harassment, sexual, violence) content so none of them were shown to any participants.
@@ A.1 Instructions for Survey Participants
A screenshot of the instructions given to participants is provided figure 4
@@ A.2 Experiemnt asking users to rate identical responses
For each prompt, the 10 sampled responses described in section 3.1 are not all unique. We ran a 20 participant experiment on Profific with 5 pairs of unique responses for 5 different prompts. We found 17 participants gave a 5 (completely equivalent) for the identical responses. Further, three participants each gave a 4, 3, and 0 to three different pairs of identical responses. This shows participants are highly likely to rate identical pairs of responses as 5 (completely equivalent). With our approach of removing highest and lowest ratings described in section 3.3, we excluded all identical pairs of responses from the user study.
@@ A.3 Supplemental Tables and Figures
We provide additional tables and figures here in support of our findings.
@@
Table 5 provides details on demographics data of participants in our study.

### v3

@@ 4 Results
Here, we first summarize the user-study data collection (4.1), the results of which were used to establish our baseline for consistency. To compare previous consistency methods to our user-rating-based method, we followed a bottom-up approach. First, we looked at how methods for estimating similarity between sentences (or LLM responses, in our case) compare to user ratings (4.2). Next, we used these similarity scores (including user-based scores) to compute consistency scores for individual responses (4.3). We also compared logit-based uncertainty-estimation methods to our user baseline. Finally, we aggregated the consistency scores within a set of responses to obtain a prompt-level consistency score (4.4). For each step of the hierarchy, we compared automated methods for estimating consistency to our user-rating-based baseline, which we believe is the closest to representing ground truth. We found that while some methods outperform others, none are very close to the user-rating-based baseline. We further found that our ensemble logit-based scores approximates the best-performing sampling method with the benefit of not needing pools of resampled responses.
@@ 4.1 User Study Overview and Demographics
We recruited 2,976 participants from Prolific, asking each to rate five unique response pairs from different prompts. About 52% of participants were female and 47% male. Participants’ ages ranged from 18 to 88 with various education, ethnicity, and income backgrounds. More details are provided in table˜6.
@@
For the user study, we removed duplicate response-pairs (including pairs with identical responses, see section˜A.2), and were left with 756 unique response-pairs, each of which was rated by five or more participants.
The prompts were on average 13.33 tokens long and the responses were 52.80 tokens long. We break down the lengths for prompts drawn from the two datasets and responses generated by the three models in table˜2.
@@ 4.2 Response-Pair Similarity
We first compare user ratings of the similarity between LLM response-pairs to semantic similarity metrics used in prior work. Semantic similarity metrics are a core component of sampling-based consistency metrics. As defined in section˜3.3, we refer to such comparison as response-pair similarities.
@@
We evaluated the semantic similarities in a response-pair using previously published methods and computed their correlation with human ratings. Each response-pair was rated by at least five participants. We calculated Krippendorff’s Alpha after removing the highest and lowest score (see section˜3.3) and found moderate agreements ($\alpha=0.72$) across raters Krippendorff (2011); Marzi et al. (2024). Comparing each of the four metrics for evaluating the similarity of response-pairs—Bert, BLEU, Rouge, and USE Zhang et al. (2020); Papineni et al. (2001); Lin (2004); Cer et al. (2018)—to participants’ ratings, we found that no single metric correlated best with how participants rated response-pairs’ similarities across all datasets and models. Notably, USE had the highest correlation coefficient with response-pairs produced by Llama, while Bert performed the best for Mistral.
Interestingly, we found the correlation between participants’ ratings and the five metrics from previous work to be higher for prompts from the CoQA dataset than the LMSYS dataset. Participants found response-pairs created from prompts from the CoQA dataset to be more similar to each other than response-pairs created from LMSYS prompts, by one level in the 6-point scale. To the best of our knowledge, Bert, BLEU, Rouge, and USE have not been used to evaluate LLM consistency with the more open-ended LMSYS dataset, which could have contributed to their lower correlation with human ratings.
Result 1: Among four existing metrics for evaluating the consistency of LLM responses based on similarity between pairs of sentences, USE best approximates human ratings in most cases. Additionally, existing metrics better approximate human ratings on CoQA than on LMSYS.
@@ 4.3 Response to Response-Set Consistency
In section˜4.2, we evaluated methods for measuring the similarity between pairs of responses. Consistency metrics, however, typically involve computing a score for each response that represents the extent to which that response is similar to or different from other responses to the same prompt. Here, we compare how human-rating-based consistency metrics perform in relation to those calculated with response-pair similarity metrics, as well as one additional metric that requires sampling multiple responses, SE Kuhn et al. (2023). As defined in section˜3.3, we call this response to response-set consistency.
@@
Motivated by reducing the cost of generating multiple responses to a prompt for consistency estimation, we further explore whether logit-based consistency-estimations methods (which have negligible cost compared to response generation) can match human-based consistency scores. For this investigation, we used 16 different logits based scores—mean, minimum, maximum, and sum for each of DLR, Entropy, Probability, and LogProbability (explained in section˜3.3)—and found none of them individually approximates the consistency of LLMs as well as the USE score (see table˜4).
@@
Figure 2: At response to response-set level, our ensemble of 16 logit-based scores is as close of an approximation of human ratings as USE.
@@
As described in section˜3.3, we used an ensemble of the 16 logit-based scores to attempt to approximate human ratings. To determine which combination of the 16 logit-based scores can create the best ensemble, we used Sequential Feature Selection (SFS)  [^7] https://scikit-learn.org/stable/modules/generated/sklearn.feature_selection.SequentialFeatureSelector.html. We ran SFS 1,600 times, performing 100 repetitions at each selection size (from using 1 logit-based score up to using all 16 scores). Across the 1,600 runs, we found maximum entropy to be most frequently selected (1,569 times) followed by sum of LogProbability (1,384 times). We provide the full list in table˜5. Within each of the 1,600 runs, we used 10-fold cross validation and evaluated the performance (i.e., Spearman $\rho$ and MSE) of our ensemble relative to human ratings. We found that using all 16 logit-based scores resulted in the highest Spearman correlation coefficient and lowest mean squared error when compared to human ratings (see figure˜5). We found our ensemble method with 16 logit-based scores had a higher correlation with human ratings than Bert, BLEU, Rouge, and SE. Our ensemble, when compared to the human ratings, performed as well as USE, with a 0.002 better MSE (figure˜2). As in section˜4.2, we found that at the response to response-set level, existing metrics correlate with human ratings better on the CoQA dataset than on LMSYS (see figure˜6).
In addition to matching the performance of the best existing metrics, our ensemble method benefits from reduced computational cost. With the same GPU, it took Bert 2,649 seconds, USE 23 seconds, Rouge 14 seconds, SE 47 seconds, and our ensembling method 0.0044 seconds to evaluate the consistency of 1,000 responses to 100 prompts. Training the ensemble with 10-fold cross-validation on 90% of the responses required an average of 0.037 seconds across 100 runs using all 16 features.
@@
Result 2: Using an ensemble of 16 logit-based scores can produce a consistency estimate that approximates human ratings as accurately as the best existing metrics. In addition to matching USE, our method greatly reduces computational cost by not requiring the generation of multiple responses.
@@ 4.4 Prompt Level Consistency
As evidenced by the difference in how existing metrics perform on the two different prompt datasets (table˜3), prompt choice is a major contributor to consistency.
For example, responses to the factual question “Where was the first modern Olympic Game?” were objectively more consistent than responses to the open-ended question “How can LLMs help users in their daily life?” As such, we next investigate how to best measure consistency for a given prompt. Specifically, we aggregate response-level consistency from section˜4.3 to obtain a single consistency score per prompt.
Figure 3: At per-prompt level, among existing metrics, our logit-based ensemble method and USE have the highest Spearman correlation coefficient with human evaluation of model-prompt consistency.
@@
Using the definition from section˜3.3, we calculate the consistency score of each prompt using participants’ ratings and metrics from previous work. We found that USE-based consistency scores best correlate with human-based consistency scores across all prompts (figure˜3). The consistency scores on the CoQA dataset correlates better to human ratings than the consistency scores on LMSYS. This is similar to the bias we observed with pair-wise comparisons in section˜4.2.
Logit-based ensemble method
Our logit-based ensemble method, described in sections˜3.3 and 4.3, performs as well as USE, when aggregated to the prompt level, in approximating human ratings of per-prompt consistency. While it did not outperform USE, our method benefits from not needing to generate a set of responses and compare responses within the set to determine LLMs’ responses consistency at the prompt level.
We further investigated the number of responses needed to train an ensembled model with logit-based scores to predict the consistency of LLM responses based on generating a single response. However, due to the relatively small number of prompts and responses in our sample, we were not able to conclude how many responses are needed for such model to accurately predict LLMs’ response consistency at the prompt level.
Result 3: Our ensemble method using logit-based scores performs as well as existing metrics in approximating human ratings of LLM prompt-level consistency. This approach enables estimation of LLM consistency without the need to generate multiple responses to a prompt.
@@ 5 Conclusion and Discussion
Using the human evaluation data, obtained via our user study, and subsequent experiments, we show that existing (automated) formulations of LLMs’ consistency are meaningfully different from our human-judgment baseline. The best existing response to response-set consistency method achieves Spearman’s $\rho=0.80$, indicating correlation but not representation. Further, prior work has suggested that correlations considered “strong” in less precise contexts Schober et al. (2018) are not as meaningful for measuring success on NLP tasks Deutsch (2022); Bavaresco et al. (2024); Shen et al. (2023). This result holds regardless of whether we examine consistency at the per-response level (section˜4.3) or at the prompt level (section˜4.4).
Though sampling-based methods come closest to the human baseline, we show that an ensemble of logit-based methods can approximate this performance ($\rho_{ensemble}=0.82$ vs $\rho_{\text{\emph{USE}}}=0.81$), creating an opportunity to avoid the sampling overhead.
We also find (section˜4.2) that the difference in human judgement-based uncertainty vs. existing metrics for LLM consistency approximation is greater for real-world prompt datasets (LMSYS) than for artificial ones (CoQA). Determining why this is the case is out of scope for our work, but we speculate that the research community’s focus on artificial testing datasets might be a contributing factor.
@@
The discrepancy between the human baseline and existing automated methods of measuring LLM consistency raises important questions: How do imperfect consistency estimation methods affect downstream tasks? How would end-users be affected if shown consistency metrics Kapoor et al. (2024)? How are models affected when consistency measurements are part of the training cycle Liu et al. (2024)? The answers are unclear and could be investigated by future work.
Finally, we urge researchers to consider the human evaluation baseline in future research on how to measure LLM consistency, as well as in research that utilizes these metrics for downstream tasks. We further recommend they (also) use real-world prompts for evaluation.
@@
Since our investigation required logit-based scores, which are often not accessible with black-box models, we used open-weights models. Such usage limits the generalizability of our findings to black-box models. Additionally, the prompts and responses in our sample are relatively short (see table˜2) despite using realistic prompts from users Zheng et al. (2023); therefore, our findings may not generalize to prompts and responses of all lengths. Lastly, user studies are expensive to conduct. While we recruited almost 3,000 users, we were only able to evaluate 100 prompts and 10 responses per prompt. As we alluded to in section˜5, more data from users are needed to establish a robust baseline for measuring LLM consistency.
@@
Our study was approved by our institution’s ethics review board. For the user study, we first provided an informed consent form to participants explaining the purpose of our survey, expected length, risks and benefits, as well as compensation. Participants who gave consent to participate proceeded to the survey. As explained in section˜3.1, responses shown to participants within the survey were filtered by the authors to remove potentially harmful (e.g., harassment, sexual, violence) content.
Distribution of Data and Artifact
Code developed for this project, anonymized data collected from participants, and analysis results are accessible at https://doi.org/10.17605/OSF.IO/T9BF5.
@@ A.1 Instructions for Survey Participants
A screenshot of the instructions given to participants is provided (figure˜4).
@@ A.2 Experiemnt asking users to rate identical responses
For each prompt, the 10 sampled responses described in section˜3.1 are not all unique. We ran a 20 participant experiment on Profific with 5 pairs of unique responses for 5 different prompts. We found 17 participants gave a 5 (completely equivalent) for the identical responses. Further, three participants each gave a 4, 3, and 0 to three different pairs of identical responses. This result show participants are highly likely to rate identical pairs of responses as a 5 (completely equivalent) so we excluded all identical pairs of responses from further study.
@@ A.3 Supplemental Tables and Figures
Logit-Based Score Name Frequency
Max. Entropy 1569
Sum. LogProb. 1384
Mean Prob. 1309
Sum. DLR 1220
Min. DLR 1123
Mean Entropy 1044
Mean DLR 984
Sum. Prob. 856
Sum. Entropy 691
Min. Entropy 678
Max. DLR 662
Min. Prob. 555
Max. LogProb. 533
Max. Prob. 414
Mean LogProb. 344
Min. LogProb. 234
Table 5: Across 1,600 SFS runs (16 features $\times$ 100 repetitions), the most frequently selected features were Mac. Entropy, Sum. LogProb., and Mean Prob.
@@
Table 6 shows participants’ demographics information.
@@
A.4 Example Prompts
Table 7 shows prompts that had large discrepancies between human rated consistency and consistency calculated with existing metric (i.e., USE).
Prompt Dataset LLM Used User Rated Consistency USE Consistency
What has Williams become with this win? CoQA Gemma 0.69 0.42
Please resume the book "a libertarian walks into a bear" LMSYS Gemma 0.50 0.28
Write to me something about Introversion and how I feel peace with small circle of people. LMSYS Llama 0.77 0.51
Write a recipe for a unique dessert. LMSYS Mistral 0.30 0.63
Table 7: Sample prompts with large differences between humans’ ratings and the best surrogate metric (i.e., USE).